# Unleashing Expert Opinion from Social Media for Stock Prediction

### v1

@@ Abstract
Index Terms:
Stock prediction, expert identification, social media analysis, graph neural networks
@@ I Introduction
Quantitative investment research has traditionally relied on volume-price data and fundamental data for stock prediction, with the goal of forecasting return ratios or price movement trends [1, 2, 3, 4, 5, 6]. In recent years, there has been a growing focus on utilizing sentiment features derived from social media as a complementary information source for stock prediction [7, 8, 9, 10, 11, 12]. Social media platforms dedicated to investment and trading discussions (e.g., StockTwits, reddit) have become increasingly influential in financial markets [13, 14]. These platforms not only reflect investor behavior and market sentiment in real time but also demonstrate significant market-moving potential. A representative example is the GameStop short squeeze [15] that originated on social media and led to unprecedented market volatility. Recognizing this potential, social media data has been actively incorporated into trading strategies by professional trading firms (e.g., WorldQuant, the top-tier hedge fund) to find predictive signals [16]. Among all the investment-related social media platforms or communities, StockTwits [17] stands out as a specialized platform designed specifically for investors to share insights on various stocks. On StockTwits, each stock has its own dedicated discussion thread, where users can post messages related to that stock. Unlike general social media platforms where sentiment of the posts typically relies on Large Language Models with potential inaccuracies, StockTwits offers a unique advantage: users can explicitly self-label their posts with “Bullish” or “Bearish” tags, providing clear and accurate sentiment indicators for specific stocks. This self-labeling mechanism eliminates the uncertainty associated with algorithmic sentiment analysis and offers more reliable sentiment signals.
@@
Although StockTwits provides valuable self-labeled sentiment data and contains a vast amount of discussions, extracting useful trading signals for either stock movement trend or return ratio prediction from the raw social media data remains challenging due to its substantial noise. To demonstrate that most social media posts are noisy and non-informative for trading predictions, we collected all StockTwits posts from 2017 to 2023 and examined the relationship between public sentiment and subsequent stock price movements. We applied a native filter and aggregation method that identifies the dominant sentiment for each stock based on a threshold of posts and the intensity of the sentiment. Specifically, if the number of posts for a given stock exceeded 30 on a given day, and one sentiment (either bullish or bearish) dominated by more than 85%, we used that sentiment to predict the stock’s trend for the following day. As shown by the solid blue line with square markers in Figure 1, the accuracy of dominant sentiment for predicting the stock’s movement trend (bullish for rise and bearish for fall) on the next day (T+1) was 47.63%, 3 days later (T+3) was 46.99%, and 7 days later (T+7) was 47.01%. These results suggest that using aggregated sentiment for stock prediction yields accuracy that is not even better than random guessing, demonstrating the limited predictive value of social media sentiment. Such underperformance indicates that the majority of social media posts contain significant noise rather than actionable trading information.
@@
In fact, social media data is inherently noisy when used as trading signals [18, 19, 20, 21, 22]. This is largely due to the open nature of these platforms, where anyone can post opinions, many of which neither influence market movements nor have predictive power. Retail investors’ discussions, for example, have limited impact on stock movements, and many of their posts on these platforms are speculative, driven by personal biases, or based on rumors [23]. These non-professional posts often lack investment insights, making them unreliable as trading signals [24]. Additionally, insider traders can deliberately manipulate social media to influence investor sentiment and, in turn, affect market trends [25]. Moreover, the prevalence of emotional and impulsive posts during market volatility further contributes to this noise, as users tend to react to price movements rather than provide predictive insights [26]. As a result, truly informative posts from expert users are rare, making it crucial to accurately identify these experts amidst the noise [20].
@@
Despite efforts in existing research to identify expert signals [18, 27, 20, 19, 28], current methods struggle to reliably distinguish true experts from users who simply appear to be experts due to random correlations with market movements. These methods often overlook the fact that true experts should be able to make precise predictions across different market regimes. Their expert identification approaches tend to favor users who post frequently and whose posts happen to be correct during specific market trends, leading to biased results. Furthermore, these methods do not address the widespread presence of bots and spam accounts in social media posts, further undermining their effectiveness in identifying genuine expert signals.
@@
To address these limitations in expert identification, we propose a novel dynamic expert tracing algorithm tailored to stock data. Our algorithm incorporates multiple key principles to ensure robust expert identification. First, it filters out noise from bots and spammers to address the data quality issue. Second, it evaluates both long-term and dynamic prediction accuracy across various market regimes to identify experts. Third, it ensures that experts focus only on a manageable number of stocks by filtering out users with numerous posts on different stocks in a single day. Furthermore, we identify both true experts and ‘inverse experts’—users who consistently make incorrect predictions. The identification of inverse experts are valuable trading signals for two main reasons. First, research in behavioral economics [29, 30] shows that several cognitive biases can lead to systematic misjudgments in market predictions, a characteristic frequently observed in inverse experts. This is particularly relevant since most traders in the market consistently lose money due to these cognitive biases, effectively making them inverse experts whose predictions can serve as reliable contrary indicators. Second, some inverse experts may emerge from deliberate market manipulation. Certain institutions or large investors strategically manipulate stock prices through social media for profit. For example, when they want to sell, they use controlled media accounts to post bullish views to encourage retail investors to buy. This allows them to offload their shares at higher prices while they sell in large volumes, causing the price to drop (see [31] for cases). Conversely, when these institutions want to buy, they post bearish views to prompt retail investors into selling. This enables them to accumulate shares at lower prices before they drive the price up. These manipulated media accounts, though deceptive in intent, can act as reliable inverse experts that provide valuable trading insights.
@@
For true experts, we can align our trading actions with their sentiment predictions (taking long positions for bullish signals and short positions for bearish ones), while for inverse experts, we can strategically take opposite positions. By applying this dual-strategy approach with our expert identification method, we predict stock price movements across different time horizons: the next day (T+1), 3 days later (T+3), and 7 days later (T+7). As shown in Figure 1, the experts identified by our expert tracing system demonstrate notably higher accuracy in predicting future stock movement trends, leading to significant improvements in expert identification over existing methods. While predicting the stock future trend is important, the primary objective of quantitative investing is not just to identify trends but to maximize excess returns by selecting stocks with the highest potential for future profits. Thus, our model aims to predict the stock return ratio for the next trading day. To achieve this, we transform the expert predictions, which serve as binary trend indicators (rise or fall), into continuous return ratios that can be treated as expert signals. Once we have used these expert signals to predict the return ratio, we can evaluate the model’s performance using various quant-related metrics that are crucial for investment strategies. These include correlation metrics that reflect the relative strength of the relationship between predicted and actual returns (e.g., information coefficient and rank information coefficient), return metrics that measure the profitability and risk-adjusted returns (e.g., annualized return and sharpe ratio), and predictive accuracy. These metrics collectively provide a comprehensive reference for evaluating investment strategies and establish a solid foundation for making informed and robust investment decisions.
@@
Though expert signals can provide strong predictions, challenges remain in their practical application. Like all methods that rely on expert predictions for trading signals, our strict expert tracing algorithm faces a common challenge of signal sparsity [19, 27, 28]. The resulting expert signals cover only about 4% of all stock-day combinations in our dataset (i.e., most stocks have no expert predictions on most trading days), making it impractical for some quantitative investment scenarios, such as cross-sectional stock selection. To mitigate this sparsity issue, and considering the high dependency between asset returns in financial markets [32, 33, 34], we aim to leverage the powerful propagation capabilities of graph neural networks to enhance the flow of information [35, 36]. Therefore, we propose a message-passing mechanism based on dual graph attention networks, termed DualGAT, to propagate expert signals across related stocks. By using various sources of relational information on stocks to construct multiple graph structures, our approach effectively broadens the coverage of expert signals throughout the dataset. Additionally, when incorporating these expert signals with other traditional financial features (e.g., price-volume and fundamental data), the combined signals consistently demonstrate enhanced predictive power in forecasting stock returns. This synergistic effect suggests that our expert signals capture distinct stock characteristics that complement, rather than merely duplicate, the information of traditional financial features. Such complementarity is particularly valuable for quantitative strategies as it diversifies the sources of information, thereby enhancing the robustness of our investment framework. The pipeline of our method can be seen in Fig 2. In summary, our contributions are as follows:
@@
  • We address the common signal sparsity challenge in expert-based trading strategies through our DualGAT model, which greatly enhances signal coverage by propagating expert predictions across related stocks.
@@
Figure 1: Comparison of stock prediction accuracy (%) between different expert identification methods and naive sentiment aggregation methods.
@@ II-A Social Media-based Stock Prediction
Stock prediction using social media data has gained significant attention in recent years due to its potential to capture market sentiment and crowd wisdom. Early research primarily focused on sentiment analysis of social media posts to predict market movements. For instance, Bollen et al. [37] analyzed Twitter mood to predict daily stock market changes, while Nguyen et al. [38] developed sentiment-based features from StockTwits messages for market prediction. These studies demonstrated that social media sentiment could serve as a valuable indicator of market trends. As deep learning advanced, especially in natural language processing, researchers began incorporating more sophisticated methods. DeepClue [39] developed a hierarchical neural network to process company-related tweets from the social media and visualize their contribution to the prediction of stock movements. Xu et al. [7] presented a market information encoder with a novel deep generative model to jointly encode tweets and price signals for stock prediction while Sawhney et al. [8] proposed MAN-SF that uses bilinear transformations to learn temporal interactions between tweet representations and historical prices. Zhang et al.[12] integrated social media sentiment as one of the model inputs in their dynamic graph attention network.
However, as mentioned in the introduction, tweets from social media are inherently noisy. Moreover, our experiments (Figure 1) show that public sentiment has limited predictive power for stock movements, which is also supported by Oliveira et al. [40] and Coyne et al. [19]. While the aforementioned studies have made progress in leveraging social media data for stock prediction, they did not effectively filter the massive amount of tweets, potentially leading to models processing largely non-informative information. This highlights the importance of identifying experts within social media platforms whose predictions can provide informative and valuable trading signals.
@@ II-B Expert Identification in Social Media
Expert identification in social media has emerged as a critical research area, with various approaches developed to address this challenge. Early studies primarily employed authority-oriented methods or topic-oriented approaches [41, 42]. Authority-oriented methods maninly focus on users’ reputation or social influence by constructing the user-to-user graph. For instance, Bouguessa et al. [43] employed an in-degree method to identify experts while Zhu et al. [44] ranked user authority in an extended category graph. Although effective, these approaches may not always match new questions with relevant expertise. In contrast, topic-oriented methods leverage latent topic modeling and linguistic analysis of users’ posts to match experts based on question content. Guo et al. and Zhou et al. [45, 46] introduced topic-based probabilistic models to question-answering activities and identify experts. Weng et al. [47] proposed TwitterRank, which identifies influential users within specific domains by performing a topic-sensitive random walk that models the transition probability from one Twitter user to another. Additionally, temporal dynamics have been recognized as crucial factors in expertise identification, as demonstrated by Chi et al. [48], who showed that dynamic LDA topic models outperform traditional static approaches in tracking evolving expertise indicators.
In the context of financial social media, expert identification primarily focuses on users’ prediction accuracy for securities. Several studies [20, 19, 18, 28, 27] have attempted to identify financial experts based on their historical prediction accuracy over a specific time period or number of posts. However, these methods fail to filter out automated bots or spammers and suffer from the problem of insufficient past data criteria when identifying expert users (see Sec.IV Expert User Identification for a detailed analysis). These limitations are effectively addressed by our dynamic expert tracing algorithm.
Figure 2: An overview of our proposed method. We first identify experts and inverse experts from social media based on their historical prediction performance across different market regimes, as shown in the left part where the prediction patterns of different types of social media users (including experts/inverse experts, bots, spammers, lucky users, and mob, see Sec.IV Expert User Identification for detailed analysis) are illustrated. Then we identify and extract these expert signals which are sparse in nature. To address the sparsity issue, we propose a dual graph attention network (DualGAT) that takes both the sparse expert signals and features obtained by the market and fundamental data as input. The DualGAT incorporates relational information among stocks from multiple aspects to propagate expert signals across related stocks, significantly expanding the coverage of expert signals, and finally outputs the daily cross-sectional return predictions for each stock.
@@ II-C Graph-based Methods in Financial Applications
Due to the phenomenon of stock momentum spillover, Graph Neural Networks (GNNs) and graph modeling have demonstrated remarkable success in financial applications by effectively modeling complex relationships between market entities [4, 6, 49, 50, 51]. Recent studies have leveraged GNN architectures to model both market relationships and temporal dynamics. Chen et al. [52] proposed a collaborative model that employs LSTM and GCN. Feng et al. [2] introduced a temporal graph convolution layer to capture stock relations in a time-sensitive manner, enabling the evolution of relational weights among connected edges over time. Hsu et al. [3] developed FinGAT, which employs graph attention networks to identify profitable stock combinations by learning latent interactions among stocks and sectors. Zhao et al. [49] proposed a bi-typed market knowledge graph approach with dual attention networks to capture momentum spillover signals between stocks and related entities such as executives through heterogeneous GNNs. Wang et al. [53] introduced HATR-I, a hierarchical adaptive temporal relational model that formulate different views of domain adjacency graphs into a unified multiplex network and use the multi-stage relational matching for stock prediction. Liu et al. [54] proposed a novel graph modeling method to explore the interactions among stock factors. Ying et al. [55] captured temporal relationships using both sequential features and attributes from stock documents through a time-aware relational attention network. Li et al. [56] transform the price series into a graph structure using chart similarity to forecast turning points in stock price.

### v2

@@ I Introduction
Quantitative investment research has traditionally relied on handcrafted features from volume-price data and fundamental indicators for stock prediction. Deep learning has significantly advanced this field by moving beyond manual feature engineering to learn complex patterns directly from raw data. Early sequence models such as LSTMs [lstm1, lstm2] and GRUs [agree_tkde, gru1] were designed to capture temporal dependencies by using the financial time series, while subsequent developments introduced Transformer-based architectures [gezici_deep_2024, tuncer_asset_2022] for enhanced sequence modeling and incorporated graph structures using GNNs to exploit inter-stock relations [feng2019temporal, fingat_tkde, tkde_Spillover]. In recent years, there has been a growing focus on utilizing sentiment features derived from social media as a complementary information source for stock prediction [acl2018stock, Deep_attentive_learning, acl2023causality, tkdd_ynamic]. Social media platforms for investment discussions have become increasingly influential in financial markets [audrino2020impact, wang2023methods]. These platforms reflect real-time investor sentiment and demonstrate significant market-moving potential, exemplified by the GameStop short squeeze [noauthor_gamestop_2024] that originated on social media and caused unprecedented volatility. Recognizing this potential, professional trading firms (e.g., WorldQuant) actively incorporate social media data into trading strategies for predictive signals [worldquant_data]. Among all investment-related social media platforms, StockTwits [stocktwits] stands out as a specialized platform for investors to share stock insights. Each stock has its dedicated discussion thread where users post related messages. Additionally, StockTwits offers a unique advantage: users can explicitly label their posts as “Bullish” or “Bearish,” providing clear sentiment indicators for specific stocks. This self-labeling mechanism eliminates algorithmic sentiment analysis uncertainty and offers more reliable sentiment signals.
@@
Although StockTwits provides valuable self-labeled sentiment data and contains a vast amount of discussions, extracting useful trading signals for either stock movement trend or return ratio prediction from the raw social media data remains challenging due to its substantial noise. To demonstrate that most social media posts are noisy and non-informative for trading predictions, we collected all StockTwits posts from 2016 to 2023 and examined the relationship between public sentiment and subsequent stock price movements. We applied a native filter and aggregation method that identifies the dominant sentiment for each stock based on a threshold of posts and the intensity of the sentiment. Specifically, if the number of posts for a given stock exceeded 30 on a given day, and one sentiment (either bullish or bearish) dominated by more than 85%, we used that sentiment to predict the stock’s trend for the following day. As shown by the solid blue line with square markers in Figure 1, the accuracy of dominant sentiment for predicting the stock’s movement trend (bullish for rise and bearish for fall) on the next day (T+1) was 47.63%, 3 days later (T+3) was 46.99%, and 7 days later (T+7) was 47.01%. These results suggest that using aggregated sentiment for stock prediction yields accuracy that is not even better than random guessing, demonstrating the limited predictive value of social media sentiment. Such underperformance indicates that the majority of social media posts contain significant noise rather than actionable trading information.
@@
In fact, social media data is inherently noisy when used as trading signals [sentiment_analysis_for, Forecating_stock_prices, aaai_expert_based, xie2020signal, schnaubelt2020separating]. This stems from the open nature of these platforms, where anyone can post opinions that often lack predictive power. Retail investors’ discussions, for example, have limited impact on stock movements, and many of their posts are speculative, driven by personal biases, or based on rumors [rumor_volatility]. These non-professional posts often lack investment insights, making them unreliable as trading signals [harvard_retail]. Additionally, there is potential for insider traders to manipulate social media to influence investor sentiment and, in turn, affect market trends [tan2023market]. Moreover, the prevalence of emotional and impulsive posts during market volatility further contributes to this noise, as users tend to react to price movements rather than provide predictive insights [ge2020beyond]. As a result, truly informative posts from expert users are rare, making it crucial to accurately identify these experts amidst the noise [aaai_expert_based].
@@
Despite efforts in existing research to identify expert signals [sentiment_analysis_for, A_user_centric, aaai_expert_based, Forecating_stock_prices, liao2014winning], current methods struggle to reliably distinguish true experts from users who simply appear to be experts due to random correlations with market movements. These methods often overlook the fact that true experts should be able to make precise predictions across different market regimes. Their expert identification approaches tend to favor users who post frequently and whose posts happen to be correct during specific market trends, leading to biased results. Furthermore, these methods do not address the widespread presence of bots and spam accounts in social media posts, further undermining their effectiveness in identifying genuine expert signals.
@@
To address these limitations in expert identification, we propose a novel dynamic expert tracing algorithm tailored to stock data. Our algorithm incorporates multiple key principles to ensure robust expert identification. First, it filters out noise from bots and spammers to address the data quality issue. Second, it evaluates both long-term and dynamic prediction accuracy across various market regimes to identify experts. Third, it ensures that experts focus only on a manageable number of stocks by filtering out users with numerous posts on different stocks in a single day. Furthermore, we identify both true experts and ‘inverse experts’—users who consistently make incorrect predictions. The identification of inverse experts are valuable trading signals for two main reasons. First, research in behavioral economics [Hilton2001ThePO, lekovic2020cognitive] shows that several cognitive biases can lead to systematic misjudgments in market predictions, a characteristic frequently observed in inverse experts. This is particularly relevant since most traders in the market consistently lose money due to these cognitive biases, effectively making them inverse experts whose predictions can serve as reliable contrary indicators. Second, inverse experts may emerge in contexts where market participants attempt to influence stock prices, particularly when certain institutions or large investors use social media to create misleading signals. For example, when they want to sell, they use controlled media accounts to post bullish views to encourage retail investors to buy. This allows them to offload their shares at higher prices while they sell in large volumes, causing the price to drop (see [sec2022] for cases). Conversely, when these institutions want to buy, they post bearish views to prompt retail investors into selling. This enables them to accumulate shares at lower prices before they drive the price up. These manipulated media accounts, though deceptive in intent, can act as reliable inverse experts that provide valuable trading insights.
For true experts, we align our trading actions with their sentiment predictions (taking long positions for bullish signals and short positions for bearish ones), while for inverse experts, we take the opposite positions. To evaluate our expert identification method, we measure predictive accuracy at different horizons: the next day (T+1), 3 days later (T+3), and 7 days later (T+7). Here, a prediction is considered correct if the direction implied by the expert’s sentiment (bullish = expected rise, bearish = expected fall, with inverse experts treated oppositely) matches the actual stock price movement over the corresponding horizon. Importantly, we identify experts separately for each time horizon, ensuring that experts for each horizon are selected based on their predictive performance within that specific timeframe and subsequently evaluated using the same corresponding horizon. When multiple experts post about the same stock on the same day, we randomly sample one expert’s sentiment per stock-day combination to ensure unbiased evaluation. As demonstrated in Figure 1, experts identified through our expert tracing system achieve substantially higher accuracy in predicting future stock movement trends across all three time horizons compared to existing expert identification methods.
@@
While predicting the stock future trend is important, the primary objective of quantitative investing is not just to identify trends but to maximize excess returns by selecting stocks with the highest potential for future profits. Thus, our model aims to predict the stock return ratio for the next trading day. To achieve this, we transform the expert predictions, which serve as binary trend indicators (rise or fall), into continuous return ratios that can be treated as expert signals. Once we use these expert signals to predict return ratios, we evaluate model performance using quantitative metrics including correlation metrics (e.g., information coefficient, rank information coefficient), return metrics (e.g., annualized return and Sharpe ratio), and predictive accuracy for comprehensive strategy evaluation.
@@
Though expert signals can provide strong predictions, challenges remain in their practical application. Like all methods that rely on expert predictions for trading signals, our strict expert tracing algorithm faces a common challenge of signal sparsity [Forecating_stock_prices, A_user_centric, liao2014winning]. The resulting expert signals cover only about 4% of all stock-day combinations in our dataset (i.e., most stocks have no expert predictions on most trading days), making it impractical for some quantitative investment scenarios, such as cross-sectional stock selection. To mitigate this sparsity issue, and considering the high dependency between asset returns in financial markets [ane2003dependence, zhu2022forecasting, zhang2023graph], we aim to leverage the powerful propagation capabilities of graph neural networks to enhance the flow of information [zhang2023adaprop, xiao2021learning]. Therefore, we propose a message-passing mechanism based on dual graph attention networks, termed DualGAT, to propagate expert signals across related stocks. By using various sources of relational information on stocks to construct multiple graph structures, our approach effectively broadens the coverage of expert signals throughout the dataset. Additionally, when incorporating these expert signals with other traditional financial features (e.g., price-volume and fundamental data), the combined signals consistently demonstrate enhanced predictive power in forecasting stock returns. This synergistic effect suggests that our expert signals capture distinct stock characteristics that complement, rather than merely duplicate, the information of traditional financial features. Such complementarity is particularly valuable for quantitative strategies as it diversifies the sources of information, thereby enhancing the robustness of our investment framework. The pipeline of our method can be seen in Fig 2. In summary, our contributions are as follows:
@@
  • We address the common signal sparsity challenge in expert-based trading strategies through our DualGAT model, which greatly enhances signal coverage by propagating expert signals across related stocks.
@@
Figure 1: Comparison of stock price movement prediction accuracy (%) achieved by experts identified through different expert identification methods versus naive sentiment aggregation methods. Accuracy is measured by how well subsequent sentiment predictions from identified experts align with actual stock price movements.
@@ II-A Social Media-based Stock Prediction
Stock prediction using social media data has gained significant attention in recent years due to its potential to capture market sentiment and crowd wisdom. Early research primarily focused on sentiment analysis of social media posts to predict market movements. For instance, Bollen et al. [bollen2011twitter] analyzed Twitter mood to predict daily stock market changes, while Nguyen et al. [nguyen2015sentiment] developed sentiment-based features from StockTwits messages for market prediction. These studies demonstrated that social media sentiment could serve as a valuable indicator of market trends. As deep learning advanced, especially in natural language processing, researchers began incorporating more sophisticated methods. DeepClue [tkde2018deepclue] developed a hierarchical neural network to process company-related tweets from the social media and visualize their contribution to the prediction of stock movements. Xu et al. [acl2018stock] presented a market information encoder with a novel deep generative model to jointly encode tweets and price signals for stock prediction while Sawhney et al. [Deep_attentive_learning] proposed MAN-SF that uses bilinear transformations to learn temporal interactions between tweet representations and historical prices. Zhang et al.[tkdd_ynamic] integrated social media sentiment as one of the model inputs in their dynamic graph attention network.
However, as mentioned in the introduction, tweets from social media are inherently noisy. Moreover, our experiments (Figure 1) show that public sentiment has limited predictive power for stock movements, which is also supported by Oliveira et al. [On_the_predictability] and Coyne et al. [Forecating_stock_prices]. While the aforementioned studies have made progress in leveraging social media data for stock prediction, they did not effectively filter the massive amount of tweets, potentially leading to models processing largely non-informative information. This highlights the importance of identifying experts within social media platforms whose predictions can provide informative and valuable trading signals.
@@ II-B Expert Identification in Social Media
Expert identification in social media has emerged as a critical research area, with various approaches developed to address this challenge. Early studies primarily employed authority-oriented methods or topic-oriented approaches [zhao_expert_2015, zhao2016expert]. Authority-oriented methods maninly focus on users’ reputation or social influence by constructing the user-to-user graph. For instance, Bouguessa et al. [bouguessa_identifying_2008] employed an in-degree method to identify experts while Zhu et al. [zhu_ranking_2014] ranked user authority in an extended category graph. Although effective, these approaches may not always match new questions with relevant expertise. In contrast, topic-oriented methods leverage latent topic modeling and linguistic analysis of users’ posts to match experts based on question content. Guo et al. and Zhou et al. [guo2008tapping, zhou2012topic] introduced topic-based probabilistic models to question-answering activities and identify experts. Weng et al. [weng2010twitterrank] proposed TwitterRank, which identifies influential users within specific domains by performing a topic-sensitive random walk that models the transition probability from one Twitter user to another. Additionally, temporal dynamics have been recognized as crucial factors in expertise identification, as demonstrated by Chi et al. [chi2018expert], who showed that dynamic LDA topic models outperform traditional static approaches in tracking evolving expertise indicators.
In the context of financial social media, expert identification primarily focuses on users’ prediction accuracy for securities. Several studies [aaai_expert_based, Forecating_stock_prices, sentiment_analysis_for, liao2014winning, A_user_centric] have attempted to identify financial experts based on their historical prediction accuracy over a specific time period or number of posts. However, these methods fail to filter out automated bots or spammers and suffer from the problem of insufficient past data criteria when identifying expert users (see Sec.IV Expert User Identification for a detailed analysis). These limitations are effectively addressed by our dynamic expert tracing algorithm.
Figure 2: An overview of our proposed method. We first identify experts and inverse experts from social media based on historical prediction performance across different market regimes, as shown in the left part illustrating prediction patterns of different user types (experts/inverse experts, bots, spammers, lucky users, mob; see Sec.IV). Then we identify and extract these expert signals which are sparse in nature. To address the sparsity issue, we propose a dual graph attention network (DualGAT) that takes both sparse expert signals and market and fundamental features as input. DualGAT incorporates relational information among stocks from multiple aspects to propagate expert signals across related stocks, expanding expert signal coverage and finally outputs the daily cross-sectional return predictions for each stock.
@@ II-C Graph-based Methods in Financial Applications
Due to the phenomenon of stock momentum spillover, Graph Neural Networks (GNNs) and graph modeling have demonstrated remarkable success in financial applications by effectively modeling complex relationships between market entities [tkde_Spillover, tkde_bitype, thgnn, samba]. Recent studies have leveraged GNN architectures to model both market relationships and temporal dynamics. Chen et al. [chen2018incorporating] proposed a collaborative model that employs LSTM and GCN. Feng et al. [feng2019temporal] introduced a temporal graph convolution layer to capture stock relations in a time-sensitive manner, enabling the evolution of relational weights among connected edges over time. Hsu et al. [fingat_tkde] developed FinGAT, which employs graph attention networks to identify profitable stock combinations by learning latent interactions among stocks and sectors. Zhao et al. [tkde_bitype] proposed a bi-typed market knowledge graph approach with dual attention networks to capture momentum spillover signals between stocks and related entities such as executives through heterogeneous GNNs. Wang et al. [wang2022hatr] introduced HATR-I, a hierarchical adaptive temporal relational model that formulate different views of domain adjacency graphs into a unified multiplex network and use the multi-stage relational matching for stock prediction. Ying et al. [ying2020time] captured temporal relationships using both sequential features and attributes from stock documents through a time-aware relational attention network. Li et al. [li2024forecasting] transform the price series into a graph structure using chart similarity to forecast turning points in stock price.